# Deep Neural Networks for Cross-Energy Particle Identification at RHIC and LHC

### v1

@@ Abstract
This work demonstrates the application of a deep neural network for particle identification in the high transverse momentum regime. A model trained on simulated Large Hadron Collider (LHC) proton-proton collisions ($\sqrt{s}=13\,\mathrm{TeV}$) is used to classify nine distinct particles using seven kinematic-level features. The model is then tested on high transverse momentum RHIC data without any transfer learning, fine-tuning, or weight adjustment. It maintains accuracy above 91% for both LHC and RHIC sets, while achieving above 96% accuracy for all RHIC sets, including the $p_{T}>7\,\mathrm{GeV}/c$ set, despite not having been trained on any RHIC data. These results indicate that the model captures the underlying physics of high-energy collisions rather than just overfitting to the training data. This work highlights the potential of simulation-trained models to be deployed in different energy domains, especially in underrepresented or data-limited settings.
keywords: Deep learning, Particle identification, Generalization, Cross-domain, High $p_{T}$, RHIC, LHC
@@ 1 Introduction
Machine learning (ML) has played an increasingly central role in data collection and analysis in particle physics. Leading studies using deep neural networks have cemented machine learning as an essential tool at CERN’s LHC [1]. In particular, deep learning has shown significant promise in particle identification tasks ranging from jet classification [2], trigger systems, and calorimetry and event reconstruction [3]. Furthermore, it has become more critical in high momentum regimes where standard particle identification methods fail [4]. This large variety of machine learning applications is mainly applied to the Large Hadron Collider (LHC) data. In comparison, a lot less attention has been given to applying such methods in the context of the Relativistic Heavy Ion Collider (RHIC), despite its unique and rich physics environment.
Proton–proton collisions at RHIC, carried out at $\sqrt{s}=200\,\mathrm{GeV}$, represent a very different kinematic regime compared to that of the LHC’s $\sqrt{s}=13\,\mathrm{TeV}$. This difference in kinematics presents a valuable opportunity to test the ability of deep ML models trained on one center-of-mass energy to generalize to another. Such cross-regime generalization is rarely explored, and most ML efforts in high-energy physics are trained and tested on the same experimental or simulated data at the same energy scale.
In this study, we tested the ability of a simple deep neural network trained exclusively on LHC data to identify nine distinct but common particles in simulated proton-proton collisions. The model was trained on seven kinematic-level features and tested on various $p_{T}>3\,\mathrm{GeV}/c$ bins. Furthermore, we tested the model’s ability to generalize to RHIC’s lower center-of-mass energy under the same high $p_{T}$ conditions as the LHC set.
To our knowledge, this work represents one of the first attempts at LHC to RHIC generalization using deep learning, highlighting both the potential and the limitations of simulation-trained models in multi-regime applications. The findings suggest that with appropriate feature selection and network architecture, similar models can be used in real experimental settings, offering a practical solution when the training data is limited, noisy, or simply not available.
@@ 2 Dataset and feature design
PYTHIA 8[5] was used to simulate the p-p collision at both $\sqrt{s}=13\,\mathrm{TeV}$ and $\sqrt{s}=200\,\mathrm{GeV}$. There are two constraints on the generated data, a pseudorapidity cut $-2<\eta<2$ which corresponds to the central region of the detector, and a transverse momentum cut $p{{}_{{}_{T}}}>$ 3 GeV/c. The hadronic decay is turned on, which makes it necessary to choose particles that are long-lived and commonly produced in proton–proton collisions. This is because the model is not designed to deal with anomalies or particles that are quite rare and produced with low multiplicities. The particles selected are $\gamma$, $\pi^{+}$, $\pi^{-}$, $K^{+}$, $K^{-}$, $K_{L}^{0}$, $p$, $e^{-}$, $n$. The training set is a combination of two generated sets with slightly different parameters to facilitate data generation and event selection. It consists of a low $p_{T}$ set that has only the pseudorapidity condition and a high $p_{T}$ set with a $p_{T}>3\,\mathrm{GeV}/c$ cut. The training set contains 6.2 million events (5 million low $p_{T}$ and 1.2 million high $p_{T}$). Table 1 shows the generation parameters in detail for the training set and Table 2 shows the selected input features along with a brief description of them.
@@
Input Feature Physical Significance Equation
Charge Electric charge of the particle in units of $e$ $q$
3 Kinematic Feature Distributions and Class Composition
The input parameters are selected based on the criterion that they should be obtainable from laboratory measurements or derivable from other measured quantities. Moreover, the detectors that inspire the feature selection should be available at both the LHC and RHIC. Based on that information, the detectors responsible for the measurements would be hadronic and electromagnetic calorimeters[6, 7], time projection chambers[8, 9], and silicon vertex detectors. These detectors are foundational and essential for particle colliders. There is an additional constraint on the parameters, which is that they must be able to be generated using the event generator Pythia 8. The distributions of the key input features for both the low- and high-$p_{T}$ sets are shown in Figure 1, illustrating the kinematic differences between the two datasets.
@@ 4 Neural Network Training
Compared to other alternatives such as tanh or sigmoid, ReLU avoids exponentiation. This makes it suitable for handling large datasets such as those from LHC simulations.Non-linearity in hidden layers in essential in capturing complex relationships between inputs and outputs.
The output layer activation function and the loss function are closely related and therefore are discussed together. The SoftMax activation function is a well suited for multi-class classification which is the goal of the network. The function converts the values it receives from the hidden layer into a probability distribution of the classes, with the sum of probabilities being equal to one. The SoftMax function takes the form: [18]
@@
For the network’s structure, each input parameter is assigned to one neuron in the input layer; therefore, the input layer has seven neurons. The number of hidden layers depends on the complexity of the task, but having two or more makes the network a deep neural network. Deep neural networks are suitable for tasks where standard methods struggle, due to their ability to capture complex patterns. The number of neurons in the hidden layer has a wide range of integer values depending on the application. However, a good starting point is having an equal number of neurons as the input layer. This network underwent many different iterations before settling on the final structure. The initial structure was 7 neurons in both hidden layers; however, to reduce overfitting, the number of neurons was reduced to 5 in the second hidden layer. Finally, the output layer uses a SoftMax activation function, where each neuron represents the probability of a class. Since there are nine particle types, the output layer contains nine neurons. Figure 2 shows a schematic representation of the network.
@@
The network was trained for 34 epochs with a batch size of 128, taking an average of 41.5 seconds per epoch and about 23.5 minutes to train. The chosen regularization method is early stopping[22] with a patience value of 8. The model continues iterating until there are no improvements in validation loss for 8 consecutive epochs. Afterwards, the model restores the weights from the epoch with the lowest validation loss. This training technique improves in generalization and reduces overfitting on future test sets.
@@
The validation consists of examples which the model hasn’t seen. It is used after each epoch to check overall how it generalizes to unseen data. The validation set is crucial in guiding and improving the performance of the model and avoiding issues such as overfitting, where the model performs well in training but fails generalizing to new examples.
@@
For the normalization of the training data, by observing the values of all the input parameters, they all fall within manageable ranges where the minimum is $-3.14$ (from the azimuthal angle). In terms of transverse momentum, most particles fall within a range of 00 to $50\,\mathrm{GeV}/c$, with a small number of outliers having $100$ to $200\,\mathrm{GeV}/c$. Naturally, for the low momentum set without the $p_{T}>3\,\mathrm{GeV}/c$ cut, most of the generated particles are soft particles having a transverse momentum value between 00 and $2\,\mathrm{GeV}/c$. Since the values for all input parameters are within close range of each other, normalization was deemed to be unnecessary.
@@
From Table 4, the model performs well on all three sets, having comparable accuracy above 99%. The lowest accuracy among the three is the training set, but that can be attributed to its large size while the validation and test sets represent only a small subset of the total data. This means the training set contains a broader range of high momentum events, which explains its slightly reduced accuracy.
@@ 6 Discussion
The key difference between the two sets is strikingly visible at the $p_{T}>7\,\mathrm{GeV}/c$ cut, where the RHIC set has much better accuracy at 96.23% compared to 91.33% of the LHC set. This difference can be explained by the training of the model and the nature of the LHC set. The model has been trained on the LHC set at $\sqrt{s}=13\,\mathrm{TeV}$, which means that the particles are much more energetic overall, with momentum ranges that are much higher than RHIC, especially at the transverse momentum range of $p_{T}>7\,\mathrm{GeV}/c$. The performance of the model on different center of mass energies can be better understood by comparing the transverse momentum distributions in Figure 4.
A dataset with the condition of $p_{T}>3\,\mathrm{GeV}/c$ includes a lot of high momentum particles, but relatively few in the $p_{T}>7\,\mathrm{GeV}/c$ region, which contributes to the drop in accuracy for LHC. For any generated set, most of the particles produced have the lowest possible energy and momentum, which corresponds to $3\,\mathrm{GeV}/c$ for the training set and $7\,\mathrm{GeV}/c$ for test set.

### v3

# Particle Identification with Deep Neural Networks Across Collision Energies in Simulated Proton–Proton Collisions

@@ Abstract
This study demonstrates a proof-of-concept application of a deep neural network for particle identification in simulated high transverse momentum proton-proton collisions, with a focus on evaluating model performance under controlled conditions. A model trained on simulated Large Hadron Collider (LHC) proton-proton collisions at $\sqrt{s}=13\,\mathrm{TeV}$ is used to classify nine particle species based on seven kinematic-level features. The model is then tested on simulated high transverse momentum Relativistic Heavy Ion Collider (RHIC) data at $\sqrt{s}=200\,\mathrm{GeV}$ without any transfer learning, fine-tuning, or weight adjustment. It maintains accuracy above 91% for both LHC and RHIC sets, while achieving above 96% accuracy for all RHIC sets, including the $p_{T}>7\,\mathrm{GeV}/c$ set, despite never being trained on any RHIC data. Analysis of per-class accuracy reveals how quantum chromodynamics (QCD) effects, such as leading particle effect and kinematic overlap at high $p_{T}$, shape the model’s performance across particle types. These results suggest that the model captures physically meaningful features of high-energy collisions, rather than simply overfitting to kinematics of the training data. This study demonstrates the potential of simulation-trained deep neural networks to remain effective across lower energy regimes within a controlled environment, and motivates further investigation in realistic settings using detector-level features and more advanced network architectures.
keywords: Deep learning, Particle identification, High transverse momentum, Proton-proton collisions, RHIC, LHC, Collider simulations
@@ 1 Introduction
Machine learning (ML), deep learning in particular, is playing an increasingly central role in data collection and analysis in modern particle physics especially at CERN’s Large Hadron Collider (LHC) [1]. Deep learning techniques have shown significant promise in particle identification tasks ranging from jet classification [2], trigger systems, to calorimetry and event reconstruction [3]. Furthermore, such approaches have become critical in high momentum regimes where standard particle identification methods fail [4]. This large variety of machine learning applications has been primarily applied to the LHC data. In contrast, significantly fewer studies have explored similar techniques at the Relativistic Heavy Ion Collider (RHIC), despite its importance in studying the quark-gluon plasma (QGP) [5] and the proton’s spin structure.
Proton–proton collisions at RHIC, carried out at $\sqrt{s}=200\,\mathrm{GeV}$ [6], represent a very different kinematic regime compared to that of the LHC’s $\sqrt{s}=13\,\mathrm{TeV}$ [4, 7]. These nuanced differences in kinematic distribution pose a meaningful challenge for evaluating the ability of deep learning models trained on one center-of-mass energy to perform when applied to another. While deep learning models show promising results in particle identification tasks, most ML efforts in high-energy physics are trained and tested on the same experimental or simulated data at the same energy scale [8, 9].
In this study, we evaluate the performance of a simple deep neural network trained exclusively on LHC proton-proton collisions to identify nine distinct but common particle types. The model was trained using seven kinematic-level features and tested on various $p_{T}>3\,\mathrm{GeV}/c$ bins. Furthermore, we assessed the model’s ability to remain effective at RHIC’s lower center-of-mass energy, using identical high-$p_{T}$ selection criteria and event counts as the LHC set.
To our knowledge, this is one of the earliest studies to explore the behavior of deep neural networks trained at LHC energies when applied to RHIC conditions in simulated proton-proton collisions. While neural networks have been used in HEP to interpolate over continuous variables like particle mass in simulated datasets [10], those approaches were primarily focused on one energy regime, rather than evaluating performance across different collision energies. This manuscript further highlights both the potential and the limitations of simulation-trained models when applied across different collider center-of-mass energies. The analysis includes per-class accuracy and misclassification patterns, providing physics-motivated insight into the model’s performance and limitations in the $p_{T}>7\,\mathrm{GeV}/c$ regime across energies. Even though the study is limited to kinematics of simulated data, the results suggest that the model can learn meaningful physical features of proton-proton collisions. The findings motivate further investigation under realistic conditions, incorporating detector-level features and advanced network architectures, with some potential directions proposed in the conclusion.
@@ 2 Dataset and Feature Design
PYTHIA 8[11] was used to simulate the p-p collision at both $\sqrt{s}=13\,\mathrm{TeV}$ and $\sqrt{s}=200\,\mathrm{GeV}$. The events were generated using PYTHIA version 8.3.07 with the Monash 2013 tune enabled by default. The default parton distribution function (PDF) used is NNPDF2.3 QCD+QED LO, which is consistent with the Monash tuning.
Two constraints are applied to the generated data, a pseudorapidity cut $-2<\eta<2$, which corresponds to the central region of the detector, and a transverse momentum cut $p{{}_{{}_{T}}}>$ 3 GeV/c. The hadronic decay is enabled, which makes it necessary to choose particles that are long-lived and commonly produced in proton–proton collisions. This is because the model is not designed to deal with anomalies or particles that are quite rare and produced with low multiplicities. The particles selected are $\gamma$, $\pi^{+}$, $\pi^{-}$, $K^{+}$, $K^{-}$, $K_{L}^{0}$, $p$, $e^{-}$, $n$.
Particles like muons, hyperons, and various short-lived particles were not selected in order to maintain charge symmetry across particle types (3 positive, 3 neutral, 3 negative) and to focus on commonly produced particles with high multiplicities within the specified kinematic cuts. Moreover, the inclusion of relatively rare particles or those requiring specialized detectors and detection methods would introduce unnecessary complexity and class imbalance without contributing to the study’s objective.
The training set is a combination of two generated sets with slightly different parameters to facilitate data generation and event selection. It consists of a low $p_{T}$ set with only the pseudorapidity condition and a high $p_{T}$ set with a $p_{T}>3\,\mathrm{GeV}/c$ cut. The training set contains 6.2 million events (5 million low $p_{T}$ and 1.2 million high $p_{T}$). Table 1 shows the selected input features along with a brief description of them and Table 2 shows the generation parameters in detail for the training set.
Input Feature Physical Significance Equation
Charge Electric charge of the particle in units of $e$ $q$
@@
3 Kinematic Distributions and Particle Composition
The input parameters are selected based on the criterion that they should be obtainable from laboratory measurements or derivable from other measured quantities. Moreover, the detectors that inspire the feature selection should be available at both the LHC and RHIC. Based on that information, the detectors responsible for the measurements would be hadronic and electromagnetic calorimeters[12, 13], time projection chambers[14, 15], and silicon vertex detectors [16, 17]. These detectors are foundational and essential for particle colliders. There is an additional constraint on the parameters, which is that they must be able to be generated using the event generator Pythia 8. The distributions of the key input features for both the low- and high-$p_{T}$ sets are shown in Figure 1, illustrating the kinematic differences between the two datasets.
@@ 4 Neural Network Training
Compared to other alternatives such as tanh or sigmoid, ReLU avoids exponentiation. This makes it suitable for handling large datasets such as those from LHC simulations. Non-linearity in hidden layers is essential in capturing complex relationships between inputs and outputs.
The output layer activation function and the loss function are closely related and therefore are discussed together. The Softmax activation function is well suited for multi-class classification which is the goal of the network. The function converts the values it receives from the hidden layer into a probability distribution of the classes, with the sum of probabilities being equal to one. The Softmax function takes the form: [26]
@@
For the network’s structure, each input parameter is assigned to one neuron in the input layer; therefore, the input layer has seven neurons. The number of hidden layers depends on the complexity of the task, but having two or more makes the network a deep neural network. Deep neural networks are suitable for tasks where standard methods struggle, due to their ability to capture complex patterns. The number of neurons in the hidden layer has a wide range of integer values depending on the application. However, a good starting point is having an equal number of neurons as the input layer. This network underwent many different iterations before settling on the final structure. The initial structure was 7 neurons in both hidden layers; however, to reduce overfitting, the number of neurons was reduced to 5 in the second hidden layer. Finally, the output layer uses a Softmax activation function, where each neuron represents the probability of a class. Since there are nine particle types, the output layer contains nine neurons.
@@
The network was trained for 34 epochs with a batch size of 128, taking an average of 41.5 seconds per epoch and about 23.5 minutes to train. The chosen regularization method is early stopping[30] with a patience value of 8. The model continues iterating until there are no improvements in validation loss for 8 consecutive epochs. Afterwards, the model restores the weights from the epoch with the lowest validation loss. This training technique improves generalization and reduces overfitting on future test sets.
@@
The validation consists of examples which the model hasn’t seen. It is used after each epoch to check overall how it generalizes to unseen data. The validation set is crucial in guiding and improving the performance of the model and avoiding issues such as overfitting, where the model performs well in training but fails to generalize to new examples.
@@
The normalization of the training data was considered but it was not applied. By observing the values of all the input parameters, they all fall within manageable ranges where the minimum is $-3.14$ (from the azimuthal angle). In terms of transverse momentum, most particles fall within a range of 00 to $50\,\mathrm{GeV}/c$, with a small number of outliers having $100$ to $200\,\mathrm{GeV}/c$. For the low momentum set without the $p_{T}>3\,\mathrm{GeV}/c$ cut, most of the generated particles have a transverse momentum value between 00 and $2\,\mathrm{GeV}/c$. Since the values for all input parameters are within close range of each other, normalization was not expected to affect the model’s performance significantly. The robust and consistently high accuracy performance of the model across different seeds in Table 4, is an indication that normalization was not a limiting factor for performance in this study.
@@
Figure 4: Confusion matrix for the test set (split from the $13\,\mathrm{TeV}$ LHC data), normalized row wise. The diagonal represents the correctly identified particles while the non diagonal elements represent misidentified particles.
Accuracy is defined as the ratio of correctly classified particles to the total number of particles in a given set:
Accuracy=Number of correctly classified particles (all classes)Total number of particles (all $\text{Accuracy}=\frac{\text{Number of correctly classified particles (all classes)}}{\text{Total number of particles (all classes)}}$ (4)
As shown in Table 4, the model performs well on all three sets, having comparable accuracy above 99%. The lowest accuracy observed is the training set, which can be attributed to its large size while the validation and test sets represent only a small subset of the total data. This means the training set contains a broader range of high momentum events, which explains its slightly reduced accuracy.
To evaluate the robustness of the model, five training runs were performed with different seeds, which correspond to different weight initializations and optimizer behavior during training. For the full-range LHC test set, the resulting accuracies varied between 99.80% and 99.88%, with a mean value of 99.86% and a sample standard deviation of 0.035. This small variation indicates that the model generalizes well to unseen data, and that its performance is stable and independent of initialization and stochastic training effects.
Figure 4 shows the identification accuracy for each particle type, as well as the misidentification pattern for each class. The per-class accuracy is defined as the ratio of correctly classified particles of a given type to the total number of actual particles of that type in the set:
[…]
This definition is equivalent to efficiency in particle identification. While purity, the fraction of correctly predicted particles of a given type relative to the total number of predicted particles of the same type, is a common metric, we focus primarily on evaluating model generalization across collision energies. For this reason, we emphasize overall classification accuracy across the two center-of-mass energies. However, purity values can be obtained directly from confusion matrices if needed.
The average per-class accuracy is 98.5%, with electrons exhibiting the lowest per-class classification accuracy. This reduced accuracy is mainly attributed to the kinematic similarity between electrons and negative pions, leading to the frequent misidentification of electrons as $\pi^{-}$. Moreover, electrons are the least represented class in the dataset, making it more difficult for the model to learn and distinguish their kinematic features from those of a more dominant class, such as pions.
@@ 6 Discussion
The key difference between the two sets is strikingly visible at the $p_{T}>7\,\mathrm{GeV}/c$ cut, where the RHIC set has much better accuracy at 96.23% compared to 91.33% of the LHC set. This difference can be explained by the training of the model and the nature of the LHC set. The model has been trained on the LHC set at $\sqrt{s}=13\,\mathrm{TeV}$, which means that the particles are much more energetic overall, with momentum ranges that are much higher than RHIC, especially at the transverse momentum range of $p_{T}>7\,\mathrm{GeV}/c$. The performance of the model on different center of mass energies can be better understood by comparing the transverse momentum distributions in Figure 5. A dataset with the condition of $p_{T}>3\,\mathrm{GeV}/c$ includes a lot of high momentum particles, but relatively few in the $p_{T}>7\,\mathrm{GeV}/c$ region, which contributes to the drop in accuracy for LHC. For any generated set, the transverse momentum distribution of particles peaks at lower values, corresponding to $3\,\mathrm{GeV}/c$ for the training set and $7\,\mathrm{GeV}/c$ for the test set.
@@
Refer to caption Refer to caption
(a) LHC $p_{T}>7\,\mathrm{GeV}/c$ (b) RHIC $p_{T}>7\,\mathrm{GeV}/c$
Figure 6: Comparison of confusion matrices between (a) LHC $\sqrt{s}=13\,\mathrm{TeV}$ and (b) RHIC $\sqrt{s}=200\,\mathrm{GeV}$ test sets with $p_{T}>7\,\mathrm{GeV}/c$ cut. Despite similar overall accuracy, variations in particle-type accuracy and relative yields reveal the differences in production dynamics between the two energy regimes.
Figure 6 further clarifies the classification process at high transverse momentum by comparing the confusion matrices of LHC and RHIC test sets with $p_{T}>7\,\mathrm{GeV}/c$ cut. For both sets, there are expected drops in accuracy for most classes but the model retains decent separation ability in this quite complex regime. However, the two critical particle types, electrons and kaons, highlight important underlying physics and model limitations.
The model fails to correctly identify electrons at high $p_{T}$ for both test sets. This is primarily a physics limitation rather than a machine learning issue. In experimental settings, electrons are distinguished from pions using detector-level features such as the shower profile in electromagnetic calorimeters, the ratio of energy deposited in calorimeters to momentum ($E$/$p$) [32], and the ionization energy loss per unit length ($\mathrm{d}E/\mathrm{d}x$) in time projection chambers [14]. Since the model only trained on kinematic features, it is realistically impossible to separate electrons from pions. Therefore, the model takes a conservative approach in trying to minimize the loss function by misidentifying electrons as the more abundant and kinematically similar class, namely the negatively charged pion ($\pi^{-}$). Since electrons represent a tiny fraction of the population for both sets, their misidentification has a negligible effect on the overall accuracy of the model.
Another noteworthy trend occurs with kaons especially in the LHC test set. The positive kaon ($K^{+}$) has the second lowest accuracy across all particles at 64.83%, with all misclassifications being positive pions ($\pi^{+}$). Even though these two particle types are significantly different in their rest mass and quark content, their kinematic signatures converge at such high transverse momentum and center-of-mass energy. In the $p_{T}>7\,\mathrm{GeV}/c$ regime, the difference in rest mass becomes negligible compared to the available energy.
Furthermore, a comparison between $K^{+}$ and $K^{-}$ in the confusion matrices reveals an asymmetric response in the model’s behavior for these two particle types. Despite the fact that both are produced with nearly equal multiplicities in the LHC set, the model exhibits superior accuracy for $K^{-}$. This is clear in the LHC confusion matrix in Figure 6 , where the $K^{-}$ identification accuracy is around 84.28% compared with 64.83% for ($K^{+}$).
This discrepancy in accuracy stems from the QCD nature of proton-proton collisions. $K^{+}$ mesons are produced more abundantly than $K^{-}$ mainly due to the leading particle effect, where hadrons sharing valence quarks with the colliding particle have a higher probability of being produced [33]. The proton’s valence quark structure ($uud$) makes it easier to produce a $K^{+}$ by combining one of the valence $u$-quarks with a $\bar{s}$-quark from the quark sea. On the other hand, the production of a $K^{-}$ meson requires both quark $\bar{u}$-quark and $s$-quark to be produced from the parton interactions. Moreover, this enhancement in $K^{+}$ production results in a broader $p_{T}$ distribution as shown in Figure 7. Negative kaons are mainly concentrated in the lowest possible $p_{T}$ bin. Across the vast majority of the higher $p_{T}$ bins, $K^{+}$ mesons dominate, which explains the significantly reduced accuracy of this particle type.
Furthermore, comparing the RHIC and LHC $K^{-}$ distributions reveals the reason behind the model’s superior performance on the $K^{-}$ mesons of the RHIC set. More than 80% of negative kaons lie in the lowest $p_{T}$ bin compared to approximately 66% for the LHC counterpart. The larger center-of-mass energy of LHC allows a larger fraction of $K^{-}$ to reach higher transverse momentum compared to RHIC.
Refer to caption Refer to caption
(a) LHC kaon distribution (b) RHIC kaon distribution
Figure 7: Fraction of total $K^{+}$ and $K^{-}$ for different $p_{T}$ bins for (a) LHC $\sqrt{s}=13\,\mathrm{TeV}$ and (b) RHIC $\sqrt{s}=200\,\mathrm{GeV}$ in the $p_{T}>7\,\mathrm{GeV}/c$ test set. The counts are normalized separately for each species in both plots to illustrate their relative $p_{T}$ distributions rather than their exact numbers, since the $K^{+}$/$K^{-}$ varies significantly between the two energies.
Refer to caption Refer to caption
(a) $K^{+}$ accuracy vs. $p_{T}$ bin centers (b) $K^{-}$ accuracy vs. $p_{T}$ bin centers
Figure 8: Per-class accuracy for $K^{+}$ and $K^{-}$ for different $p_{T}$ bins for both LHC and RHIC in the $p_{T}>7\,\mathrm{GeV}/c$ test set. Accuracy is consistently higher for RHIC data within the same $p_{T}$ bins compared to LHC, which highlights fundamental differences in production dynamics between the two collision energies. A decrease in accuracy is observed for both $K^{+}$ and $K^{-}$ with increasing values of transverse momentum with the trend being more pronounced in the LHC case.
Figure 8 shows per-class accuracy of $K^{+}$ and $K^{-}$ as a function of transverse momentum. The model performs significantly better on RHIC kaons compared to LHC across all the selected momentum bins, with the gap being wider at higher momenta. This result supports the interpretation that the model’s classification ability is influenced by the underlying QCD production mechanisms, which differ meaningfully between center-of-mass energies. The lower available energy for RHIC restricts the kinematic reach of all particle species, reducing the kinematic overlap among particle types. These results suggest that the model’s behavior is driven by physical features of high-energy collisions rather than memorization of training set kinematics.